OpenStudy (anonymous):

help pleas! suppose the population of deer in a state is 10280 and is growing 1% each year. predict the population after 5 years. a: 51400 b:15556 c:10804 d:514

OpenStudy (anonymous):

compounded growth (end) = (start) (1 + rate/year)^(number of years) end = (10280) (1.01)^5

OpenStudy (anonymous):

ok ill just type that on the calculator?

OpenStudy (anonymous):

yes then use your common sense 1% of 10280 is 102.80 = 103 approx. five years of about 100 added willadd about 500, so answer needs to be close to 10280 + 500 approx. Pretty evident from options.

OpenStudy (anonymous):

ok I only got c... is that right?

OpenStudy (anonymous):

Quite right!

OpenStudy (anonymous):

can you still help me? ps/ wh cant I keep answering I have to keep logging in and out?

OpenStudy (anonymous):

Sorry that I do not know this system very well. Yuo should not have to log out/in/out each time.

OpenStudy (anonymous):

a 4000.00 principal earns 5% intrest compounded annually after 4 years what is the balance in the account? a:5000000.00 b:4862.03 c:4600.00 d:20250.00 can you help me with this?

OpenStudy (anonymous):

amount = (principal) (1 + .05)^4 amount = 1.216 Principal approximation is (Principal) (1 + 4 x 0.05) = 1.20 Principal